# Benjamin S. Manning MIT John J. Horton

### v3

@@ Appendix D Additional Figures
Notes: This table shows the results of the statistical tests comparing the strategic AI agents to the other models for $\varepsilon=0.05$. The first column shows the comparison model. The second presents $e^{\bar{\Lambda}_{S}}$ with bootstrap confidence intervals comparing the strategic AI agents to the other models. The third and fourth columns present p-values for the Wilcoxon rank-sum test and random-sign permutation test, respectively. The fifth column presents the proportion of games for which the strategic AI agents is the best predictor with its 95% Clopper-Pearson interval. Significance Indicator: ∗∗∗p<<0.001, ∗∗p<<0.01, p<<0.05.
@@
Notes: This table shows the results of the statistical tests comparing the strategic AI agents to the other models for $\varepsilon=0.1$. The first column shows the comparison model. The second presents $e^{\bar{\Lambda}_{S}}$ with bootstrap confidence intervals comparing the strategic AI agents to the other models. The third and fourth columns present p-values for the Wilcoxon rank-sum test and random-sign permutation test, respectively. The fifth column presents the proportion of games for which the strategic AI agents is the best predictor with its 95% Clopper-Pearson interval. Significance Indicator: ∗∗∗p<<0.001, ∗∗p<<0.01, p<<0.05.
@@
Notes: This table shows the results of the statistical tests comparing the strategic AI agents to the other models for $\varepsilon=0.2$. The first column shows the comparison model. The second presents $e^{\bar{\Lambda}_{S}}$ with bootstrap confidence intervals comparing the strategic AI agents to the other models. The third and fourth columns present p-values for the Wilcoxon rank-sum test and random-sign permutation test, respectively. The fifth column presents the proportion of games for which the strategic AI agents is the best predictor with its 95% Clopper-Pearson interval. Significance Indicator: ∗∗∗p<<0.001, ∗∗p<<0.01, p<<0.05.
@@
Notes: This table shows the results of the statistical tests comparing the strategic AI agents to the other models for $\varepsilon=0.3$. The first column shows the comparison model. The second presents $e^{\bar{\Lambda}_{S}}$ with bootstrap confidence intervals comparing the strategic AI agents to the other models. The third and fourth columns present p-values for the Wilcoxon rank-sum test and random-sign permutation test, respectively. The fifth column presents the proportion of games for which the strategic AI agents is the best predictor with its 95% Clopper-Pearson interval. Significance Indicator: ∗∗∗p<<0.001, ∗∗p<<0.01, p<<0.05.
Table D7: Summary statistics comparing human responses with model-predicted supports $\varepsilon=0$.
Optimized AI Baseline AI HS Nash Eq.
% Humans Choose Max Prob. Strategy 24.3 16.8 30.4
% Humans Choose Top 3 Prob. Strategy 52.9 39.1 49.6
% Humans Choose Pos. Prob. Strategy 94.3 81.9 46.4
% Games Any Human Chooses Pos. Prob. Strategy 99.3 93.7 74.7
% Games All Humans Choose Pos. Prob. Strategy 86.3 65.3 17.7
Notes: Each column reports regression estimates within a subgroup defined by the points rule or bonus rule. We report Huber-White robust standard errors. The reference categories are the “normal” points rule and the “coordinate low” bonus rule.
@@
Log-Likelihood Ratio
Baseline AI HS Nash Eq.
(1) (2)
Normal -1 (Pts) -0.878∗∗∗ (0.260) 0.419 (0.348)
Normal -2 (Pts) -1.148∗∗∗ (0.256) 0.290 (0.340)
Normal +1 (Pts) -0.692∗∗ (0.267) 0.510 (0.340)
Normal +2 (Pts) -0.680∗∗ (0.258) 0.250 (0.343)
Two Less Max Costless (Pts) -0.899∗∗ (0.308) -1.147∗∗ (0.384)
Equal (Bonus) -0.439 (0.414) -0.341 (0.521)
Gap Absolute (Bonus) -0.705 (0.420) 0.151 (0.492)
Gap Higher (Bonus) -1.573∗∗∗ (0.392) 0.591 (0.528)
Gap Lower (Bonus) -1.130∗∗ (0.415) -0.139 (0.473)
Less Upper (Bonus) 0.340 (0.438) 1.270 (0.537)
More Than (Bonus) -0.916 (0.414) 1.880∗∗∗ (0.511)
Sum Even (Bonus) -1.190∗∗ (0.400) 1.160 (0.547)
Sum Odd (Bonus) -0.942 (0.414) 0.442 (0.535)
Sum Upper (Bonus) 0.726 (0.461) 0.957 (0.574)
Unequal (Bonus) -0.545 (0.429) 0.045 (0.514)
Constant 3.202∗∗∗ (0.388) 1.859∗∗∗ (0.448)
Observations 1,477 1,477
R2 0.070 0.039
Adjusted R2 0.061 0.029
Notes: Each column reports regression estimates within a subgroup defined by the points rule or bonus rule. We report Huber-White robust standard errors. The reference categories are the “normal” points rule and the “coordinate low” bonus rule. Significance Indicator: ∗∗∗p<<0.001, ∗∗p<<0.01, p<<0.05.
@@
Log-Likelihood Ratio
Baseline AI HS Nash Eq.
(1) (2)
Normal -1 (Pts) -0.724∗∗∗ (0.219) 0.387 (0.294)
Normal -2 (Pts) -0.927∗∗∗ (0.216) 0.270 (0.287)
Normal +1 (Pts) -0.586∗∗ (0.226) 0.444 (0.288)
Normal +2 (Pts) -0.567∗∗ (0.220) 0.214 (0.292)
Two Less Max Costless (Pts) -0.824∗∗ (0.255) -0.826 (0.328)
Equal (Bonus) -0.104 (0.351) -0.493 (0.441)
Gap Absolute (Bonus) -0.383 (0.352) 0.123 (0.416)
Gap Higher (Bonus) -1.086∗∗∗ (0.325) 0.275 (0.453)
Gap Lower (Bonus) -0.711 (0.344) -0.097 (0.401)
Less Upper (Bonus) 0.300 (0.364) 0.930 (0.458)
More Than (Bonus) -0.622 (0.345) 1.408∗∗ (0.432)
Sum Even (Bonus) -0.745 (0.337) 0.781 (0.468)
Sum Odd (Bonus) -0.516 (0.351) 0.267 (0.452)
Sum Upper (Bonus) 0.593 (0.379) 0.637 (0.492)
Unequal (Bonus) -0.288 (0.359) 0.105 (0.431)
Constant 2.510∗∗∗ (0.322) 1.181∗∗ (0.383)
Observations 1,477 1,477
R2 0.056 0.033
Adjusted R2 0.046 0.023
Notes:Each column reports regression estimates within a subgroup defined by the points rule or bonus rule. We report Huber-White robust standard errors. The reference categories are the “normal” points rule and the “coordinate low” bonus rule. Significance Indicator: ∗∗∗p<<0.001, ∗∗p<<0.01, p<<0.05.
@@
Log-Likelihood Ratio
Baseline AI HS Nash Eq.
(1) (2)
Normal -1 (Pts) -0.553∗∗ (0.174) 0.349 (0.239)
Normal -2 (Pts) -0.686∗∗∗ (0.174) 0.247 (0.234)
Normal +1 (Pts) -0.467 (0.182) 0.373 (0.234)
Normal +2 (Pts) -0.441 (0.178) 0.182 (0.239)
Two Less Max Costless (Pts) -0.727∗∗∗ (0.201) -0.512 (0.271)
Equal (Bonus) 0.180 (0.284) -0.629 (0.361)
Gap Absolute (Bonus) -0.079 (0.280) 0.124 (0.342)
Gap Higher (Bonus) -0.631 (0.256) 0.005 (0.377)
Gap Lower (Bonus) -0.315 (0.270) -0.011 (0.329)
Less Upper (Bonus) 0.236 (0.288) 0.599 (0.376)
More Than (Bonus) -0.365 (0.272) 0.964∗∗ (0.352)
Sum Even (Bonus) -0.354 (0.269) 0.400 (0.387)
Sum Odd (Bonus) -0.154 (0.282) 0.111 (0.368)
Sum Upper (Bonus) 0.455 (0.295) 0.338 (0.407)
Unequal (Bonus) -0.071 (0.284) 0.175 (0.347)
Constant 1.791∗∗∗ (0.253) 0.536 (0.317)
Observations 1,477 1,477
R2 0.042 0.027
Adjusted R2 0.032 0.017
Notes: Each column reports regression estimates within a subgroup defined by the points rule or bonus rule. We report Huber-White robust standard errors. The reference categories are the “normal” points rule and the “coordinate low” bonus rule. Significance Indicator: ∗∗∗p<<0.001, ∗∗p<<0.01, p<<0.05.
@@
Log-Likelihood Ratio
Baseline AI HS Nash Eq.
(1) (2)
Normal -1 (Pts) -0.439∗∗ (0.146) 0.321 (0.205)
Normal -2 (Pts) -0.531∗∗∗ (0.146) 0.230 (0.201)
Normal +1 (Pts) -0.386 (0.153) 0.328 (0.200)
Normal +2 (Pts) -0.355 (0.150) 0.165 (0.206)
Two Less Max Costless (Pts) -0.648∗∗∗ (0.166) -0.335 (0.235)
Equal (Bonus) 0.308 (0.239) -0.691 (0.312)
Gap Absolute (Bonus) 0.089 (0.235) 0.146 (0.296)
Gap Higher (Bonus) -0.382 (0.211) -0.116 (0.329)
Gap Lower (Bonus) -0.098 (0.224) 0.069 (0.284)
Less Upper (Bonus) 0.189 (0.241) 0.416 (0.325)
More Than (Bonus) -0.228 (0.225) 0.731 (0.303)
Sum Even (Bonus) -0.157 (0.225) 0.184 (0.337)
Sum Odd (Bonus) 0.019 (0.237) 0.041 (0.316)
Sum Upper (Bonus) 0.377 (0.243) 0.182 (0.354)
Unequal (Bonus) 0.035 (0.236) 0.224 (0.296)
Constant 1.347∗∗∗ (0.210) 0.179 (0.276)
Observations 1,477 1,477
R2 0.035 0.026
Adjusted R2 0.025 0.016
Notes: Each column reports regression estimates within a subgroup defined by the points rule or bonus rule. We report Huber-White robust standard errors. The reference categories are the “normal” points rule and the “coordinate low” bonus rule. Significance Indicator: ∗∗∗p<<0.001, ∗∗p<<0.01, p<<0.05.

### v4

@@ D.3 Robustness checks for the family of pre-committed games in Section 4
Notes: This table shows the results of the statistical tests comparing the strategic AI agents to the other models for $\varepsilon=0.05$. No smoothing is applied to the cognitive hierarchy model. The first column shows the comparison model. The second presents $e^{\bar{\Lambda}_{S}}$ with bootstrap confidence intervals comparing the strategic AI agents to the other models. The third and fourth columns present p-values for the Wilcoxon signed-rank test and random-sign permutation test, respectively. The fifth column presents the proportion of games for which the strategic AI agents is the best predictor with its 95% Clopper-Pearson interval. Significance Indicator: ∗∗∗p<<0.001, ∗∗p<<0.01, p<<0.05.
@@
Notes: This table shows the results of the statistical tests comparing the strategic AI agents to the other models for $\varepsilon=0.1$. No smoothing is applied to the cognitive hierarchy model. The first column shows the comparison model. The second presents $e^{\bar{\Lambda}_{S}}$ with bootstrap confidence intervals comparing the strategic AI agents to the other models. The third and fourth columns present p-values for the Wilcoxon signed-rank test and random-sign permutation test, respectively. The fifth column presents the proportion of games for which the strategic AI agents is the best predictor with its 95% Clopper-Pearson interval. Significance Indicator: ∗∗∗p<<0.001, ∗∗p<<0.01, p<<0.05.
@@
Notes: This table shows the results of the statistical tests comparing the strategic AI agents to the other models for $\varepsilon=0.2$. No smoothing is applied to the cognitive hierarchy model. The first column shows the comparison model. The second presents $e^{\bar{\Lambda}_{S}}$ with bootstrap confidence intervals comparing the strategic AI agents to the other models. The third and fourth columns present p-values for the Wilcoxon signed-rank test and random-sign permutation test, respectively. The fifth column presents the proportion of games for which the strategic AI agents is the best predictor with its 95% Clopper-Pearson interval. Significance Indicator: ∗∗∗p<<0.001, ∗∗p<<0.01, p<<0.05.
@@
Notes: This table shows the results of the statistical tests comparing the strategic AI agents to the other models for $\varepsilon=0.3$. No smoothing is applied to the cognitive hierarchy model. The first column shows the comparison model. The second presents $e^{\bar{\Lambda}_{S}}$ with bootstrap confidence intervals comparing the strategic AI agents to the other models. The third and fourth columns present p-values for the Wilcoxon signed-rank test and random-sign permutation test, respectively. The fifth column presents the proportion of games for which the strategic AI agents is the best predictor with its 95% Clopper-Pearson interval. Significance Indicator: ∗∗∗p<<0.001, ∗∗p<<0.01, p<<0.05.
Figure D4: Relative predictive accuracy of the strategic AI agent vs other models for different game types ($\varepsilon=0.05$)
Notes: This figure shows the proportion of games for which strategic AI agents are the best predictor of initial play, separated by bonus rule (top panel) and points rule (bottom panel). All distributions except the cognitive hierarchy model are smoothed with $\varepsilon=0.05$. The vertical dashed line represents equal performance (50-50 split). Green indicates that strategic AI agent significantly outperforms the reference model in more than 50% of games, red indicates significantly worse performance in more than 50% of games, and grey indicates no significant difference. Error bars show 95% Clopper-Pearson confidence intervals.
Figure D5: Relative predictive accuracy of the strategic AI agent vs other models for different game types ($\varepsilon=0.1$)
Notes: This figure shows the proportion of games for which strategic AI agents are the best predictor of initial play, separated by bonus rule (top panel) and points rule (bottom panel). All distributions except the cognitive hierarchy model are smoothed with $\varepsilon=0.1$. The vertical dashed line represents equal performance (50-50 split). Green indicates that strategic AI agent significantly outperforms the reference model in more than 50% of games, red indicates significantly worse performance in more than 50% of games, and grey indicates no significant difference. Error bars show 95% Clopper-Pearson confidence intervals.
Figure D6: Relative predictive accuracy of the strategic AI agent vs other models for different game types ($\varepsilon=0.2$)
Notes: This figure shows the proportion of games for which strategic AI agents are the best predictor of initial play, separated by bonus rule (top panel) and points rule (bottom panel). All distributions except the cognitive hierarchy model are smoothed with $\varepsilon=0.2$. The vertical dashed line represents equal performance (50-50 split). Green indicates that strategic AI agent significantly outperforms the reference model in more than 50% of games, red indicates significantly worse performance in more than 50% of games, and grey indicates no significant difference. Error bars show 95% Clopper-Pearson confidence intervals.
Figure D7: Relative predictive accuracy of the strategic AI agent vs other models for different game types ($\varepsilon=0.3$)
Notes: This figure shows the proportion of games for which strategic AI agents are the best predictor of initial play, separated by bonus rule (top panel) and points rule (bottom panel). All distributions except the cognitive hierarchy model are smoothed with $\varepsilon=0.3$. The vertical dashed line represents equal performance (50-50 split). Green indicates that strategic AI agent significantly outperforms the reference model in more than 50% of games, red indicates significantly worse performance in more than 50% of games, and grey indicates no significant difference. Error bars show 95% Clopper-Pearson confidence intervals.
@@
Log-Likelihood Ratio (Strategic AI / {…})
{Baseline AI} {HS Nash Eq.} {Cognitive Hierarchy} {Random Pure} {Uniform}
(1) (2) (3) (4) (5)
Bonus Size 0.022 -0.046∗∗ 0.033 0.042 0.028∗∗
(0.013) (0.017) (0.014) (0.020) (0.010)
Lower Bound 0.035 -0.020 -0.033 -0.012 -0.019
(0.014) (0.018) (0.015) (0.021) (0.011)
Action Space Size 0.098∗∗∗ 0.064∗∗ -0.009 0.165∗∗∗ 0.035∗∗
(0.017) (0.022) (0.017) (0.025) (0.012)
Gap -0.012 0.211 0.078 -0.126 -0.052
(0.070) (0.092) (0.076) (0.106) (0.052)
Constant 0.139 1.960∗∗∗ 0.782 5.406∗∗∗ -0.048
(0.334) (0.441) (0.353) (0.510) (0.250)
Observations 1,477 1,477 1,477 1,477 1,477
R2 0.029 0.015 0.008 0.031 0.013
Notes: Each column reports regression estimates where the dependent variable is the log-likelihood ratio of the strategic AI agent to the other models. The independent variables are the game features from Table 3 (excluding the bonus rule and points rule). We report heteroskedasticity-robust standard errors. Significance Indicator: ∗∗∗p<<0.001, ∗∗p<<0.01, p<<0.05.
@@
Log-Likelihood Ratio (Strategic AI / {…})
{Baseline AI} {HS Nash Eq.} {Cognitive Hierarchy} {Random Pure} {Uniform}
(1) (2) (3) (4) (5)
Bonus Size 0.020 -0.032 0.031 0.039 0.026∗∗
(0.011) (0.014) (0.014) (0.017) (0.009)
Lower Bound 0.028 -0.018 -0.031 -0.011 -0.017
(0.011) (0.015) (0.014) (0.018) (0.009)
Action Space Size 0.088∗∗∗ 0.051∗∗ 0.001 0.150∗∗∗ 0.045∗∗∗
(0.014) (0.019) (0.017) (0.021) (0.011)
Gap -0.011 0.172 0.089 -0.103 -0.041
(0.058) (0.078) (0.074) (0.089) (0.047)
Constant 0.038 1.176∗∗ 0.767 3.917∗∗∗ -0.063
(0.279) (0.370) (0.345) (0.428) (0.223)
Observations 1,477 1,477 1,477 1,477 1,477
R2 0.032 0.013 0.008 0.036 0.019
Notes: Each column reports regression estimates where the dependent variable is the log-likelihood ratio of the strategic AI agent to the other models. The independent variables are the game features from Table 3 (excluding the bonus rule and points rule). We report heteroskedasticity-robust standard errors. Significance Indicator: ∗∗∗p<<0.001, ∗∗p<<0.01, p<<0.05.
@@
Log-Likelihood Ratio (Strategic AI / {…})
{Baseline AI} {HS Nash Eq.} {Cognitive Hierarchy} {Random Pure} {Uniform}
(1) (2) (3) (4) (5)
Bonus Size 0.017 -0.019 0.029 0.034 0.024∗∗
(0.009) (0.012) (0.013) (0.014) (0.008)
Lower Bound 0.020 -0.015 -0.029 -0.009 -0.015
(0.009) (0.012) (0.014) (0.014) (0.008)
Action Space Size 0.075∗∗∗ 0.036 0.007 0.131∗∗∗ 0.051∗∗∗
(0.011) (0.015) (0.017) (0.017) (0.009)
Gap -0.009 0.133 0.100 -0.079 -0.030
(0.047) (0.063) (0.073) (0.072) (0.040)
Constant -0.073 0.478 0.785 2.492∗∗∗ -0.045
(0.221) (0.299) (0.338) (0.343) (0.191)
Observations 1,477 1,477 1,477 1,477 1,477
R2 0.036 0.009 0.008 0.042 0.028
Notes: Each column reports regression estimates where the dependent variable is the log-likelihood ratio of the strategic AI agent to the other models. The independent variables are the game features from Table 3 (excluding the bonus rule and points rule). We report heteroskedasticity-robust standard errors. Significance Indicator: ∗∗∗p<<0.001, ∗∗p<<0.01, p<<0.05.
@@
Log-Likelihood Ratio (Strategic AI / {…})
{Baseline AI} {HS Nash Eq.} {Cognitive Hierarchy} {Random Pure} {Uniform}
(1) (2) (3) (4) (5)
Bonus Size 0.015 -0.012 0.027 0.031∗∗ 0.022∗∗
(0.007) (0.010) (0.013) (0.012) (0.007)
Lower Bound 0.016 -0.013 -0.027 -0.008 -0.013
(0.008) (0.011) (0.014) (0.012) (0.007)
Action Space Size 0.066∗∗∗ 0.024 0.008 0.115∗∗∗ 0.052∗∗∗
(0.009) (0.013) (0.017) (0.014) (0.008)
Gap -0.007 0.108 0.107 -0.063 -0.023
(0.039) (0.054) (0.072) (0.060) (0.035)
Constant -0.132 0.134 0.809 1.702∗∗∗ -0.021
(0.184) (0.256) (0.334) (0.290) (0.167)
Observations 1,477 1,477 1,477 1,477 1,477
R2 0.038 0.007 0.007 0.046 0.034
Notes: Each column reports regression estimates where the dependent variable is the log-likelihood ratio of the strategic AI agent to the other models. The independent variables are the game features from Table 3 (excluding the bonus rule and points rule). We report heteroskedasticity-robust standard errors. Significance Indicator: ∗∗∗p<<0.001, ∗∗p<<0.01, p<<0.05.